OpenStudy (anonymous):

Megan works a 38-hour week and for any extra time she is paid at time and a half. When she worked a 45-hour week she received $582. What would she earn for a week in which she worked 40 hours?

OpenStudy (e.cociuba):

let x be what she is paid if work time =< 38 hours therefore, 38x + 7(1.5x) = 582 = 38x+10.5x = 582 = 48.5x = 582 x=12 hence, she is paide $12 /h if her work time =< 38 but she works 40 hours so it's: 38(12) + 2(12 x 1.5) = 492 hence, she earns $492 for a 40 hour week

OpenStudy (anonymous):

@e.cociuba Thank you so much! You're a life saver. :)

OpenStudy (e.cociuba):

np:)